# From Memorization to Creativity: LLM as a Designer of Novel Neural-Architectures

### v1

@@ 6 Limitations
Despite the observed improvements, several limitations and threats to validity remain. First, the study is restricted to a single dataset and task, namely CIFAR-10 image classification with a fixed input resolution. This choice enables controlled analysis, but it remains unclear how well the observed trends transfer to substantially different datasets (e.g., higher-resolution images or non-visual domains), input modalities, or task formulations such as segmentation or detection. The induced architectural prior may therefore be partly specialized to the CIFAR-10 setting.
Second, first-epoch validation accuracy is used as the sole performance signal. While early-epoch metrics and learning-curve extrapolation can correlate with eventual performance (domhan2015lcextrapation; ru2020speedy), the relationship is imperfect and may favor architectures that learn quickly early but plateau later. In addition, longer-horizon training of the best discovered architectures is not evaluated here, and direct comparisons to human-designed baselines or classical NAS methods under matched computational budgets are not included. As a result, the conclusions primarily concern the relative evolution of the generator under a fixed proxy evaluation protocol rather than absolute state-of-the-art performance.
Third, the refinement strategy and selection policy are intentionally simple. LoRA adaptation is performed with fixed hyperparameters across cycles, and the acceptance threshold (40% first-epoch accuracy) is held constant after selection. Although this design isolates the effect of dataset evolution, it is not necessarily optimal; the observed plateauing behavior in later cycles suggests that alternative curricula or additional regularization could be beneficial.
Finally, novelty is defined using MinHash–Jaccard similarity over token-level shingles of source code and therefore operates at the text level rather than on explicit computation graphs. Different implementations can encode functionally similar architectures, while functionally equivalent models may be treated as novel if expressed in sufficiently different styles. Although the text-level filter removes obvious near-duplicates and trivial edits, it does not capture deeper semantic equivalence between architectures.
@@ 7 Ethical Considerations
This study employs a code-oriented large language model (LLM), specifically DeepSeek-Coder-7B-Instruct-v1.5, within a closed, controlled, and iterative architecture-synthesis framework. The research is strictly methodological in nature and focuses on neural architecture search. It does not involve human participants, personal data, or the deployment of an end-user-facing system.
Data Usage and Privacy.
Security and Integrity of Generated Code.
Transparency and Reproducibility.
@@ A.2 Operational novelty filtering during sampling
Appendix B Additional Results
B.1 First-epoch accuracy trends
@@
Figure 4: Best, mean, and median first-epoch accuracy per cycle on CIFAR–10 (see Section 4).
B.2 Proportion exceeding the 40% first-epoch accuracy threshold
@@
Figure 5: Proportion of models with first-epoch accuracy $\geq 40\%$ per cycle (see Section 4).
@@
After cycle 18, both best and mean first-epoch accuracies plateau and fluctuate within a narrow band, and the fraction of models above 40% stabilizes. This saturation suggests diminishing returns from continued self-training and is consistent with known failure modes of iterative self-training pipelines, where selection bias and distributional narrowing can limit further gains when the training set becomes increasingly dominated by model-generated samples (Amini2022SelfTraining; Shumailov2024ModelCollapse).

### v2

@@ 6 Limitations
Despite the observed improvements, several limitations remain and should be weighed when interpreting the results.
Benchmark scope. Although we evaluate on three datasets (CIFAR-10, CIFAR-100, SVHN), all three are low-resolution ($32{\times}32$) image classification benchmarks. It remains unclear how well the observed trends transfer to higher-resolution images, non-visual domains, or tasks such as segmentation or detection.
Proxy signal. First-epoch validation accuracy serves as the sole performance signal. While supported by recent work on early-discarding strategies [11] and contextualized by baseline calibration in Section 4.1, validating the proxy ranking under full training schedules remains an important next step (Suppl. §E).
Scope of comparisons. This study characterizes the generator’s evolution rather than proposing a complete NAS pipeline; fully-trained accuracy comparisons against established NAS methods are therefore not included. Table 1 positions our proxy-level results alongside reported fully-trained accuracies from related LLM-based methods to provide context.
Qualitative coverage. A preliminary qualitative inspection of representative architectures (Suppl. §H) confirms structural diversity spanning compact single-block designs to deep residual networks. A more systematic analysis across a larger sample would further strengthen this evidence.
Fine-tuning rigidity. LoRA adaptation is performed with fixed hyperparameters, and the acceptance threshold is held constant; the observed plateauing in later cycles suggests that alternative curricula or additional regularization could be beneficial.
Text-level novelty definition. Novelty is defined over token-level shingles of source code rather than on explicit computation graphs, meaning functionally equivalent models expressed in different styles may be treated as novel, while structurally distinct models with similar token patterns may not.
@@ 7 Ethical Considerations
This study employs DeepSeek-Coder-7B-Instruct-v1.5 within a closed, controlled, iterative architecture-synthesis framework and is strictly methodological in nature. It does not involve human participants, personal data, or end-user-facing deployment. The initial training corpus (LEMUR) consists exclusively of source code and technical metadata containing no personally identifiable information. Any real-world deployment of LLM-generated code would necessitate comprehensive security reviews. Full details on data usage, security considerations, and reproducibility measures are provided in Suppl. §M.
Acknowledgments. This work was partially supported by the Alexander von Humboldt Foundation.
Supplementary Material
From Memorization to Creativity:
LLM as a Designer of Novel Neural Architectures
Waleed Khalid        Dmitry Ignatov        Radu Timofte
Computer Vision Lab, CAIDAS & IFI, University of Würzburg, Germany
This document provides supplementary details for the main paper. Section references of the form “Section X” refer to the main paper; supplementary sections are labelled A–L.
@@ A.2 Operational Novelty Filtering During Sampling
Appendix B LEMUR Corpus Deduplication and Conversion
The reduction from 109,913 raw LEMUR records to 1,065 unique records reflects the high degree of text-level redundancy in the raw export: many entries are minor variants of the same architecture differing only in hyperparameter values, comments, or variable names. The MinHash–LSH deduplication retains one representative per near-duplicate cluster, preserving architectural diversity while removing redundant instances that would otherwise bias fine-tuning toward overrepresented families. The 216 records dropped during chat-format conversion were excluded due to missing required interface components (e.g., absent forward method signatures, incompatible constructor arguments, or malformed metadata fields) that would have produced syntactically invalid training targets; including them would introduce noise into the supervised signal.
While this reduction inevitably narrows coverage relative to the full LEMUR corpus, the retained 849 records span a diverse range of architecture families—including standard convolutional networks, residual architectures, depthwise-separable and lightweight designs, and multi-branch topologies—across multiple image-classification task types as catalogued in the LEMUR metadata [13], providing sufficient diversity to initialize the generator without strong family-level bias. We acknowledge that architectures well-represented in dropped or deduplicated clusters may be underrepresented in the initial prior, and that this constitutes a potential source of coverage bias that future work could address by applying more permissive conversion or by supplementing with additional corpora.
Appendix C Dual-Task Training Pair Construction
For each converted record, two training examples are constructed from the same assistant code but with user messages describing the task as either MNIST ($\times28$ grayscale) or CIFAR-10 ($\times32$ RGB) classification. This duplication serves two purposes: first, it encourages the LLM to associate the same architectural pattern with multiple input specifications, promoting generalization of learned design priors across related image-classification tasks rather than overfitting to a single dataset description; second, it doubles the number of gradient updates over each architectural pattern during fine-tuning, which stabilizes learning on a corpus of fewer than 900 unique examples without introducing new architectural information. The code is identical across both variants by design: the intent is not to generate dataset-specific architectures but to train the generator to respond to task descriptions as a flexible interface while producing architectures that respect the shared structural constraints. This duplication does not inflate the number of unique architectural patterns in the training set, which remains 849.
Appendix D Stopping Criterion
The synthesis loop was run for 22 cycles, a number determined by the convergence behavior of the valid generation rate and mean first-epoch accuracy rather than by an arbitrary practical cutoff. Specifically, both metrics exhibit clear plateau behavior after approximately cycle 18 (Section 4 of the main paper), consistent with the diminishing-returns regime documented in iterative self-training literature [35, 3]. Four additional cycles (19–22) were run beyond this apparent plateau to confirm stabilization rather than temporary fluctuation, yielding a total of 22 cycles.
In practice, we recommend the following data-driven stopping criterion for future applications of this loop: terminate when the improvement in mean first-epoch accuracy across three consecutive cycles falls below a threshold $\epsilon$ (e.g., $\epsilon=0.5$ percentage points), and the proportion of models exceeding the accuracy threshold has stabilized within its Wilson confidence interval across those same cycles. Under this criterion applied retrospectively to our data, termination would have occurred at cycle 18–19, yielding results within one percentage point of the 22-cycle endpoint at approximately 18% lower computational cost.
Appendix E Justification of First-Epoch Accuracy as Performance Proxy
We adopt true validation accuracy after the first training epoch as our primary performance signal, rather than relying on zero-shot NAS proxies [11]. Although training-free indicators have been shown to correlate with fully-trained accuracy [28, 7, 27], they remain indirect. Even the strongest reported Spearman rank-correlation coefficients ($\rho\approx 0.50$$0.82$) on standard benchmarks such as NAS-Bench-101 and NAS-Bench-201 correspond to a coefficient of determination of at most $R^{2}\approx 0.67$, leaving substantial variance unexplained [1, 38]. Moreover, recent work has established that early discarding after a single epoch constitutes an effective and principled strategy in automated model selection [11], providing theoretical grounding for our proxy choice. Unlike zero-cost indicators, first-epoch accuracy is a direct measurement of how well an architecture supports learning on the target task and dataset, making it both interpretable and actionable as a selection criterion within the iterative synthesis loop.
Appendix F Detailed Ablation Analysis
This section expands on the ablation results summarized in Table 4 of the main paper.
Removing the novelty filter.
The novelty filter is disabled while the 40% accuracy threshold and LoRA fine-tuning remain unchanged. Under this condition, any architecture that compiles, trains for one epoch, and exceeds the 40% first-epoch accuracy threshold becomes eligible for inclusion in the training corpus, even if its code closely matches models already present. Across cycles, the valid generation rate and first-epoch accuracy trends remain qualitatively similar to the full method, and the generator still transitions into a regime where a large fraction of valid models exceed the 40% threshold. However, the composition of the training corpus changes substantially: without the novelty constraint, the corpus accumulates repeated motifs and near-duplicate architectures, and the number of genuinely distinct code patterns added per cycle is noticeably lower than in the full method. Generated models therefore retain acceptable accuracy but exhibit reduced code-level diversity, indicating that the novelty filter is central for sustaining exploration of new regions of the design space rather than repeatedly reinforcing a narrow family of designs.
Removing the accuracy threshold.
The MinHash-based novelty filter is retained, but the 40% first-epoch accuracy threshold is removed. In this variant, all architectures that (i) compile and train for one epoch and (ii) are non-duplicates with respect to the accepted set and earlier generations are added to the training corpus regardless of early-epoch performance. Improvements in valid generation rate are still observed as the LLM is exposed to more executable code, and the generator continues to produce architectures with reasonable first-epoch accuracy. Nonetheless, the shift in the overall accuracy distribution is clearly weaker than in the full method. The training set contains a broader mixture of low- and high-performing architectures, and low-accuracy yet valid models are regularly promoted into the corpus. Consequently, the fraction of models exceeding the 40% accuracy threshold increases more slowly and stabilizes at a lower level than when performance filtering is applied. This outcome indicates that novelty alone does not suffice to steer refinement toward rapidly learning architectures under the early-epoch proxy; performance-based selection is required to align corpus growth with empirical learning behavior.
Removing iterative fine-tuning.
The iterative refinement mechanism is removed entirely. The base LLM is fine-tuned once on the initial LEMUR-derived training split, and architectures are generated from this fixed model without adding self-generated examples back into the training corpus. The evaluation protocol is unchanged, but the feedback loop is disabled. Under this non-iterative baseline, the valid generation rate and first-epoch accuracies match the behavior observed in early cycles of the full loop and do not exhibit the progressive improvements obtained under cycle-wise updates. The proportion of models surpassing the 40% accuracy threshold remains substantially below the levels achieved in later cycles, and the distribution of architectures does not shift into the regime where high-accuracy models dominate. Although code-level novel architectures can still be produced, these discoveries do not influence future generations, leaving the generator effectively static.
Appendix G Additional Results
G.1 First-Epoch Accuracy Trends
@@
Figure 4: Best, mean, and median first-epoch accuracy per cycle on CIFAR–10 (see Section 4 of the main paper).
G.2 Proportion Exceeding the 40% First-Epoch Accuracy Threshold
@@
Figure 5: Proportion of models with first-epoch accuracy $\geq 40\%$ per cycle (see Section 4 of the main paper).
@@
After cycle 18, both best and mean first-epoch accuracies plateau and fluctuate within a narrow band, and the fraction of models above 40% stabilizes. This saturation suggests diminishing returns from continued self-training and is consistent with known failure modes of iterative self-training pipelines, where selection bias and distributional narrowing can limit further gains when the training set becomes increasingly dominated by model-generated samples [3, 35].
G.3 On Pre-training Bias and Architectural Exploration
A natural concern is whether the LLM’s generative behavior remains constrained by architectural patterns encountered during pre-training, effectively limiting exploration to known motifs such as VGG-style stacks or ResNet-like blocks. We argue this concern, while legitimate, is precisely what the iterative fine-tuning loop is designed to address, and that our results provide empirical evidence against severe constraint.
First, the base model (DeepSeek-Coder-7B) was pre-trained on general code corpora rather than curated neural architecture datasets; architectural patterns constitute a small and heterogeneous fraction of its pre-training distribution. The dominant inductive bias is therefore toward syntactically valid Python and PyTorch idioms, not toward any particular architectural family.
Second, and more directly, the MinHash–Jaccard novelty filter explicitly rejects candidates whose token-level shingle representation matches any architecture in the LEMUR-derived training corpus above threshold $\tau=0.90$. Since the LEMUR corpus itself spans a broad range of architectural families, any architecture that passes this filter is by construction dissimilar to known patterns at the code level. The 455 architectures admitted over 22 cycles on CIFAR-10 therefore represent structures that are not near-duplicates of the 109,913 LEMUR records—including the original pre-deduplicated corpus—providing a lower bound on the degree to which the generator escapes its initial prior.
Third, the monotonic increase in both the number of novel models admitted per cycle (1 in cycle 1 to 40 in cycle 17) and the diversity of structural fingerprints suggests that the feedback loop actively expands rather than contracts the explored region of design space. The cross-dataset experiments reinforce this conclusion: the loop admits novel architectures at comparable rates on CIFAR-100 and SVHN, demonstrating that exploration is not confined to CIFAR-10-specific motifs.
Appendix H Qualitative Analysis of Discovered Architectures
To complement the quantitative metrics reported in the main paper, we examine four representative architectures produced by the generator, selected to illustrate the structural diversity of the output distribution. Table 8 summarizes each architecture.
Arch. 1-ep. Acc. Architecture summary Key features
[…]
Ultra-compact single-block, parameterized num. classes
[…]
VGG-style, Kaiming init, global avg pooling
Table 8: Representative generated architectures spanning different design families. All architectures were generated within the 22-cycle synthesis loop on CIFAR-10 and passed the MinHash–Jaccard novelty filter (i.e., each is structurally dissimilar to both the LEMUR seed corpus and other generated models at the token-shingle level).
Structural diversity.
The four examples illustrate that the generator produces architectures spanning a substantial range of structural complexity. At one extreme, Architecture C is an ultra-compact design consisting of a single convolutional feature block followed immediately by global average pooling and a linear classifier—a minimal architecture that nonetheless achieves 58.1% first-epoch accuracy. At the other extreme, Architecture A is a deep multi-stage network with four make_layer stages, each containing multiple Conv–BN–ReLU blocks with progressive channel widening (64 $\to$ 512) and spatial downsampling, achieving 62.8%.
Emergence of residual connections.
Architecture B is particularly notable: the generator independently produces a Block class implementing identity shortcuts with conditional downsampling, structurally analogous to a ResNet basic block. The forward pass explicitly computes out += identity with a separate downsampling branch when spatial dimensions or channel counts change. This residual pattern, combined with a deep 3-stage layout (3/4/6 blocks) and a two-layer classifier with dropout regularization, represents a sophisticated design that passes the novelty filter—indicating it is not a near-duplicate of any architecture in the LEMUR corpus despite its conceptual similarity to ResNet. This suggests the generator is capable of re-deriving well-known effective patterns through the iterative feedback loop rather than simply copying them from training data.
Variation in design philosophy.
Architecture D adopts a VGG-style philosophy—no skip connections, uniform dual-Conv–BN–ReLU blocks per stage—but replaces the traditional flattened fully-connected classifier with global average pooling followed by a single linear layer, a more modern and parameter-efficient choice. The contrast between Architectures B and D (residual vs. plain, complex vs. uniform) illustrates that the generator maintains meaningful design diversity even among models with similar accuracy levels (61.4% vs. 58.0%).
Summary.
The examples above demonstrate that the generator produces architectures exhibiting both combinatorial novelty (different arrangements and depths of known building blocks) and operational diversity (e.g., the emergence of residual identity shortcuts in Architecture B, dropout regularization strategies, and the choice between global average pooling and fully-connected classifiers). These variations go beyond superficial code-level edits: they reflect distinct computational graphs with different gradient-flow properties and capacity–efficiency trade-offs, all synthesized within the constrained design space and confirmed as structurally novel by the MinHash–Jaccard filter.
Appendix I Prompt and API Contract Specification
Each generated model is required to implement a fixed API contract: a class Net(nn.Module) with methods __init__, forward, train_setup, and learn, together with a module-level function supported_hyperparameters() returning {"lr", "momentum"}. Prompts require import torch and import torch.nn as nn and instruct the model to output a single nn.Module definition without data loading or training loops. No pretrained weights or external feature extractors are permitted; only standard convolutional, pooling, normalization, and activation operations are allowed. An implicit edge-friendly latency target accompanies the 500,000500{,}000-parameter budget. The system message casts the model as an expert PyTorch architecture designer optimizing for first-epoch accuracy under these constraints; the user message specifies the dataset name, input/output shapes, parameter budget, and the API contract listed above. This prompt template is held fixed across all 22 cycles.
Appendix J Full Hyperparameter Specifications
LoRA fine-tuning.
Fine-tuning uses DeepSeek-Coder-7B-Instruct-v1.5 with LoRA applied to attention projections (q/k/v/o) and MLP projections (up_proj, down_proj, gate_proj) in all 24 Transformer layers (0–23). LoRA hyperparameters are fixed across cycles: rank r=32r{=}32, LoRA $\alpha{=}32$, dropout 0.05. The task is causal language modeling over chat-format prompt–response pairs. Each cycle runs for 5 epochs with learning rate $1\times 10^{-5}$, per-device batch size 1, gradient accumulation 4 (effective batch size 4), paged AdamW in 8-bit, cosine schedule with 20 warmup steps, weight decay 0.01, max gradient norm 1.0, and bfloat16 mixed precision.
Decoding configuration.
Generation uses a chat interface with fixed decoding across all cycles: temperature 0.20, top-$k$ 50, nucleus p=0.9p{=}0.9, do_sample=True, and maximum new tokens 2,048 (padded with EOS). These settings are deliberately held constant so that changes in generator behavior are attributable solely to the evolving training corpus.
Appendix K Statistical Confidence Intervals
Let $N_{\mathrm{gen}}^{(c)}$ denote the number of candidate architectures sampled in cycle $c$, and $N_{\mathrm{valid}}^{(c)}$ the subset that compile and train successfully. The valid generation rate is $p_{\mathrm{valid}}^{(c)}=N_{\mathrm{valid}}^{(c)}/N_{\mathrm{gen}}^{(c)}$.
For all valid models in cycle $c$, let $\mathcal{A}^{(c)}=\{A(m):m\in\mathcal{M}_{\mathrm{valid}}^{(c)}\}$. For the sample mean $\bar{A}^{(c)}$ and standard deviation $s^{(c)}$ computed over $n_{c}=|\mathcal{A}^{(c)}|$, we report $t$-based 95% confidence intervals:
$\bar{A}^{(c)}\pm t_{0.975,\,n_{c}-1}\frac{s^{(c)}}{\sqrt{n_{c}}}.$ (2)
For any proportion $\hat{p}=k/n$ (e.g., $p_{\mathrm{valid}}^{(c)}$ or the proportion with $A(m)\geq\tau$), we report Wilson score confidence intervals [39].
Appendix L Computational Complexity & Accessibility
To evaluate the practical accessibility of our framework, we benchmarked the computational requirements on consumer-grade hardware. Utilizing a single NVIDIA GeForce RTX 4090 (24GB), candidate generation averages 0.7–1.5 GPU-minutes per sample, while the low-fidelity proxy evaluation (1-epoch training) requires 1–5 GPU-minutes per valid model. Parameter-efficient fine-tuning via LoRA (rank 32, 5 epochs) over the terminal corpus of 2,153 pairs is completed in approximately 2–4 hours. We estimate the cumulative compute budget for the entire 22-cycle evolution at 90–266 GPU-hours, demonstrating that NNGPT enables sophisticated neural architecture design without requiring enterprise-scale clusters.
Appendix M Ethical Considerations (Extended)
Data Usage and Privacy.
Security and Integrity of Generated Code.
Transparency and Reproducibility.